OpenStudy (anonymous):

5 ----- t=-9 solve 13

OpenStudy (anonymous):

you want to solve for "t"?

OpenStudy (anonymous):

You have 5/13 * t on the left and -9 on the right Do you know how to get t by itself?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

if I has 2x = 1, and wanted to get x alone I would divide both sides by 2. 2x/2 = x, 1/2 = 1/2 x = 1/2 In your case, what do you need to divide by?

OpenStudy (anonymous):

divide both sides by 5/13...

OpenStudy (anonymous):

you divide by what you want to get rid of

OpenStudy (anonymous):

you following?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so what is on the right side?

OpenStudy (anonymous):

-9/ (5/13)

OpenStudy (anonymous):

you can invert 5/13 and multiply: -9*13/5